OpenStudy (anonymous):

how many ways are there to arrange the word ' success' such that all three 's' are separated?

OpenStudy (anonymous):

do you have the answer

OpenStudy (anonymous):

is it 720?

OpenStudy (akshay_budhkar):

final answer is not important.. what is the method you used and?

OpenStudy (anonymous):

i wasn't sure about the answer so was just asking if the answer was known. i worked it out by multiplying 3!x5!

OpenStudy (akshay_budhkar):

why? 3!*5!?

OpenStudy (anonymous):

cause there are four other letters minus the 3 "s"

OpenStudy (anonymous):

i don't think I'm right though

OpenStudy (anonymous):

s-s-s-- s-s--s- s-s---s s--s-s- s--s--s s---s-s -s-s-s- -s-s--s -s--s-s --s-s-s

OpenStudy (akshay_budhkar):

whats this?

OpenStudy (anonymous):

ten possibilities to arrange the s's, the other letters don't matter yet.

OpenStudy (anonymous):

so the answer is 10*4!

OpenStudy (anonymous):

4! for the other letters.

OpenStudy (anonymous):

hi sorry but i've no answer it's my test question today..

OpenStudy (anonymous):

my working's (4!x (4x3x2) )/3!x2!

OpenStudy (zarkon):

\[{5\choose2}\cdot\frac{4!}{2!}\]

OpenStudy (akshay_budhkar):

Zarkon enters Zarkon says answer is ... Zarkon leaves lol

OpenStudy (zarkon):

as shown above by Thomas9 there are 10 ways to arrange the blanks and s's which is \[{4+1\choose3}\] the 4blanks and 3 s's the 4! is the arrangement of the 4 non s's. But we have 2 c's that can be arranged 2! ways. therefore we have \[4!2!\] ways to arrange the four non s's so the total number of ways to arrange is thus \[{4+1\choose3}⋅4!2!=10⋅4⋅3=120\]